OpenStudy (anonymous):

Simplify the expression. cosine x ^2 + sinx ^ 2 / cot x ^ 2 - csc x ^2

OpenStudy (anonymous):

\[\cos ^{2}x + \sin ^{2}x = 1 \] and \[\cot ^{2} x= \frac{ \cos ^{2} x }{ \sin ^{2}x }\] and \[\csc ^{2}x = \frac{ 1 }{ six ^{2}x }\] knowing this can you simplify the expression?

OpenStudy (anonymous):

Yeah for the bottom part of it would I just cross multiply

OpenStudy (anonymous):

not cross multiply just use some algebra and substitution.

OpenStudy (loser66):

is it not that cot^2 - csc ^2 =-1? therefore it is = -1?